### v2

@@ 5 Properties of the Gale demand correspondence
Let $M\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sup_{y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b\log u(y)-\left\langle\mathbf{p},y\right\rangle$, which implies the existence of a sequence of allocation $\{y^{(k)}\}_{k=1}^{\infty}$ such that $b\log u(y^{(k)})-\left\langle\mathbf{p},y^{(k)}\right\rangle$ converges to $M$. We show the existence of an allocation $y^{*}$ satisfying $b\log u(y^{*})-\left\langle\mathbf{p},y^{*}\right\rangle=M$.
According to Lemma 5.1, there exists a sequence of $\bar{y}^{(k)}$ such that $b\log u(\bar{y}^{(k)})-\left\langle\mathbf{p},\bar{y}^{(k)}\right\rangle$ converges to $M$, with $0\leq\left\langle\mathbf{p},\bar{y}^{(k)}\right\rangle\leq b$ for all $k$. Thus, we can select a subsequence $\bar{y}^{(i_{k})}$ such that $b\log u(\bar{y}^{(i_{k})})-\left\langle\mathbf{p},\bar{y}^{(i_{k})}\right\rangle$ converges to $M$ and $\left\langle\mathbf{p},\bar{y}^{(i_{k})}\right\rangle$ converges to $b^{\prime}$ for some $b^{\prime}\in[0,b]$. Consequently, $b\log u(x^{(i_{k})})$ converges to $M+b^{\prime}$ as $\left\langle\mathbf{p},x^{(i_{k})}\right\rangle$ converges to $b^{\prime}$. By Assumption 1, there exists an allocation $y^{*}$ such that $b\log u(y^{*})=M+b^{\prime}$ and $\left\langle\mathbf{p},y^{*}\right\rangle\leq b^{\prime}$. This completes the proof. ∎
We can use the subgradient Karush–Kuhn–Tucker (KKT) conditions to characterize solutions in the Gale demand correspondence. We refer the reader to Lagrangian duality and KKT conditions of nondifferentiable functions to Chapter 3 in Ruszczyński (2011).
Lemma 5.3.
For a utility function $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, $b>0$ and $\mathbf{q}\in\operatorname{pdom}(u)$, let $y\in\mathcal{GD}^{u}(\mathbf{q},b)$. Then, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(y)$ such that
$g_{j}\leq{q_{j}}\cdot\frac{u(y)}{b}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }y_{j}>0\,,$
Conversely, if such a $\mathbf{g}\in\partial^{\wedge}u(y)$ exists then $y\in\mathcal{GD}^{u}(\mathbf{q},b)$.
We are ready to prove Lemma 3.1
See 3.1
@@
By Lemma 5.3, there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(y)$ such that $g_{j}\leq q_{j}u(y)/b$ with equality whenever $y_{i}>0$. Therefore,
$u(x)$ $\leq u(y)+\left\langle\mathbf{g},x-y\right\rangle\leq u(y)+\frac{u(y)}{b}\left\langle\mathbf{q},x-y\right\rangle$
$\leq u(y)\left(1+\frac{\left\langle\mathbf{q},x\right\rangle}{b}\right)\leq u(y)\left(1+\frac{\left\langle\mathbf{p},x\right\rangle}{b}\right)\leq 2u(y)\,.$
5.1 Nash welfare as approximate competitive equilibrium
We now derive Theorem 1.2, showing that Nash welfare maximizing allocations can be seen as approximate competitive equilibria.
Theorem 5.4 (Restatement of Theorem 1.2).
Consider a Fisher market instance, and let $\mathbf{y}$ be an allocation maximizing Nash welfare along with Gale prices $\mathbf{q}$. Then, $(\mathbf{y},\mathbf{q})$ forms a $2$-demand-approximate competitive equilibrium. Additionally, when all $b_{i}$’s are identical, the allocation is $2$-envy-free.
Proof.
By Lemma 5.1, all agents satisfy the budget constraints. Additionally, the market clears by the KKT conditions for (EG). For the approximate competitive equilibrium, we only need to show $u_{i}(y_{i})\geq\frac{1}{2}\max_{\left\langle\mathbf{q},x_{i}\right\rangle\leq b_{i}}u_{i}(x_{i})$ for all agents.
Note that $y_{i}$ maximizes $b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle$. By Lemma 5.3, there is a supergradient $g_{i}$ such that $g_{ij}\leq q_{j}u_{i}(y_{i})/b_{i}$ with equality whenever $y_{ij}>0$. Then, for any $x_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ such that $\left\langle\mathbf{q},x_{i}\right\rangle\leq b_{i}$,
$u_{i}(x_{i})\leq u_{i}(y_{i})+\left\langle g_{i},x_{i}-y_{i}\right\rangle\leq u_{i}(y_{i})+\frac{u_{i}(y_{i})}{b_{i}}\left\langle\mathbf{q},x_{i}-y_{i}\right\rangle\leq\left(2-\frac{\left\langle\mathbf{q},y_{i}\right\rangle}{b_{i}}\right)u_{i}(y_{i})\leq 2u_{i}(y_{i})\,.$
This proves that $(\mathbf{y},\mathbf{q})$ forms a $2$-approximate competitive equilibrium. Additionally, allocation is $2$-envy-free follows directly from Lemma 2.10. ∎
6 Efficiency of Nash welfare for Gale-substitutes
In this section, we prove Theorem 1.3. We show that for $\Sigma$-Gale subsitutes utility functions, given a competitive equilibrium $(\mathbf{x},\mathbf{p})$, there are Gale prices $\mathbf{q}^{\star}$ such that $\mathbf{q}^{\star}\leq\mathbf{p}$. The statement then follows by Lemma 3.1. We start by re-stating and proving Lemma 3.2.
See 3.2
Proof.
According to the KKT conditions, there exists $\lambda\geq 0$ and supergradient $\mathbf{g}\in\partial^{\wedge}u(x)$ such that
and $\left\langle\mathbf{p},x\right\rangle=b$ if $\lambda>0$.
Let us first consider the case $\lambda=0$. Then, $g_{j}\leq 0$ with equality if $x_{j}>0$, implying that $x=\arg\max_{x\in\mathbb{R}^{\mathcal{M}}}u(x)$. Consequently, $u(.)$ is satiable; therefore, $\mathbf{0}\in\operatorname{pdom}(u)$ and $x\in\mathcal{GD}^{u}(\mathbf{0},b)$. Now, ($\Sigma$) guarantees the existence of $y\in\mathcal{GD}^{u}(\mathbf{p},b)$, $y\leq x$. Clearly, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ in this case.
Next, assume $\lambda>0$. Let us define
$\mathbf{q}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\frac{\lambda b}{u(x)}\cdot\mathbf{p}\,.$
Note that $\mathbf{q}\leq\mathbf{p}$. This follows since $\lambda b=\lambda\left\langle\mathbf{p},x\right\rangle=\left\langle\mathbf{g},x\right\rangle\leq u(x)$, where the final inequality is by concavity and $u(\mathbf{0})=0$. Thus,
$g_{j}\leq\frac{u(x)}{b}q_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
According to Lemma 5.3, $x\in\mathcal{GD}^{u}(\mathbf{q},b)$.
Note that $\mathcal{GD}^{u}(\mathbf{q},b)=\mathcal{GD}^{u}(\mathbf{p},u(x)/\lambda)$. Utilizing Definition 2.14(ii) for $x\in\mathcal{GD}^{u}(\mathbf{p},u(x)/\lambda)$ and $b\leq u(x)/\lambda$, there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$ as required. ∎
We next show that for any competitive equilibrium $(\mathbf{x},\mathbf{p})$, and any prices $\mathbf{q}\leq\mathbf{p}$, there exist an aggregate Gale demand such that $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ whenever $q_{j}=p_{j}$.
Lemma 6.1.
Consider a Fisher market instance such that $u_{i}(\cdot)$ is $\Sigma$-Gale-substitute utilities for all agents. For any competitive equilibrium $(\mathbf{x},\mathbf{p})$, for any $\mathbf{q}\in\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ such that $\mathbf{q}\leq\mathbf{p}$, there exists $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$ and $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ such that $q_{j}=p_{j}$.
Proof.
By Lemma 3.2, there exists $\bar{y}_{i}\in\mathcal{GD}_{i}(\mathbf{p},b_{i})$ for every $i\in{\mathcal{A}}$ such that $\bar{y}_{i}\leq x_{i}$. Using Definition 2.14(i), there exists $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ and $y_{ij}\leq\bar{y}_{ij}$ if $q_{j}=p_{j}$. Thus, if $q_{j}=p_{j}$ then $\sum_{i\in{\mathcal{A}}}y_{ij}\leq\sum_{i\in{\mathcal{A}}}\bar{y}_{ij}\leq\sum_{i\in{\mathcal{A}}}x_{ij}\leq 1$. ∎
@@ 6 Efficiency of Nash welfare for Gale-substitutes
$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{y_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q},y_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$
The following properties are easy to verify.
Lemma 6.2.
@@
    Given $\mathbf{q}\in\operatorname{dom}(\varphi)$, let $y_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$. Then, $\mathbf{1}-\sum_{i\in{\mathcal{A}}}y_{i}\in\partial^{\vee}\varphi(\mathbf{q})$.
@@
Let $(\mathbf{x},\mathbf{p})$ be a competitive equilibrium solution. Consider the dual (5) of the Eisenberg-Gale program (EG), i.e.,
@@
We show that this has an optimal solution with $\mathbf{q}\leq\mathbf{p}$. The theorem then follows by Lemma 3.1. Let us take any optimal $q^{\star}\in\operatorname{dom}(\varphi)$; we are done if $\mathbf{q}^{\star}\leq\mathbf{p}$. Otherwise, let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,\mathbf{q}^{\star}_{j}>p_{j}\}$, and let us define $\bar{\mathbf{q}}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
$\bar{q}_{j}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{cases}p_{j}\,,&\mbox{if }j\in J\,,\\ q^{\star}_{j}\,,&\mbox{if }j\notin J\,.\\ \end{cases}$
@@
Let us pick an allocation $\mathbf{y}$ for the prices $\bar{\mathbf{q}}$ as guaranteed by Lemma 6.1: $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ with $\bar{q}_{j}=p_{j}$; in particular, for all $j\in J$. According to Lemma 6.2(ii), we get a subgradient $\bar{\mathbf{g}}\in\partial^{\vee}\varphi(\bar{\mathbf{q}})$ defined as $\bar{\mathbf{g}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{1}-\sum_{i\in{\mathcal{A}}}y_{i}$; this has the property $\bar{g}_{j}\geq 0$ for all $j\in J$. We then have
@@
It is left to show that $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$. Let $K\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,\mathbf{q}^{\star}_{j}=0<p_{j}\}$. Clearly, $K\cap J=\emptyset$. If $K=\emptyset$, then Lemma 5.2(iii) already implies $\bar{\mathbf{q}}\in\operatorname{pdom}(u)$, since $\mathbf{p}\geq\bar{\mathbf{q}}$ and $\operatorname{supp}(\bar{\mathbf{q}})=\operatorname{supp}(\mathbf{p})$. Otherwise, for $k=1,2,\ldots$, let us define $\mathbf{q}^{(k)}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
$q_{j}^{(k)}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\begin{cases}p_{j}\,,&\mbox{if }j\in J\,,\\ \frac{p_{j}}{k}\,,&\mbox{if }j\in K\,,\\ q^{\star}_{j}\,,&\mbox{if }j\notin J\cup K\,.\\ \end{cases}$
@@
We have $\mathbf{q}^{(k)}\in\operatorname{pdom}(u)$ again by Lemma 5.2(iii). As in the first part, we take a subgradient $\mathbf{g}^{(k)}$ as guaranteed by Lemma 6.2(ii), with the properties that $\bar{g}_{j}\geq 0$ for all $j\in J$ and also $\mathbf{g}^{(k)}\leq\mathbf{1}$. We can write
@@
$\varphi(\mathbf{q}^{\star})\geq\varphi(\mathbf{q}^{(k)})+\left\langle\mathbf{g}^{(k)},\mathbf{q}^{\star}-\mathbf{q}^{(k)}\right\rangle=\varphi(\mathbf{q}^{(k)})+\sum_{j\in J}\bar{g}_{j}(q^{\star}_{j}-q^{(k)}_{j})-\sum_{j\in K}\bar{g}_{j}q^{(k)}_{j}\geq\varphi(\mathbf{q}^{(k)})-\frac{\|p\|_{\infty}}{k}\,.$
@@
Here, we use in the first equality that $q^{\star}_{j}=q^{(k)}_{j}$ for $j\notin J\cup K$, and $q^{\star}_{j}=0$ for $j\in K$. The last inequality uses $\bar{g}_{j}\leq 1$ and the definition $q^{(k)}_{j}=p_{j}/(k+1)$. Thus, Lemma 5.2(v) is applicable with $M=\varphi(\mathbf{q}^{\star})+\|p\|_{\infty}$, using that for each $j\in{\mathcal{M}}$, $\max_{y_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b_{i}\log u_{i}(y_{i})-\left\langle\mathbf{q}^{(k)},y_{i}\right\rangle\leq\varphi(\mathbf{q}^{(k)})$. This shows that $\bar{\mathbf{q}}\in\operatorname{pdom}(u_{i})$ for all $i\in{\mathcal{M}}$, and therefore $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{M}}}\operatorname{pdom}(u_{i})$. ∎
7 Separable utilities are $\Sigma$-Gale-substitutes

### v3

@@ Appendix B Efficiency of Nash welfare for Gale-substitutes
According to the KKT conditions, there exists $\lambda\geq 0$ and supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ such that
and $\left\langle\mathbf{p},\mathbf{x}\right\rangle=b$ if $\lambda>0$.
Let us first consider the case $\lambda=0$. Then, $g_{j}\leq 0$ with equality if $x_{j}>0$, implying that $\mathbf{x}=\arg\max_{\mathbf{x}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{x})$. Consequently, $u(.)$ is satiable; therefore, $\mathbf{0}\in\operatorname{pdom}(u)$ and $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{0},b)$. Now, ($\Sigma$) guarantees the existence of $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$, $\mathbf{y}\leq\mathbf{x}$. Clearly, $\operatorname{pdom}(u)=\mathbb{R}_{\geq 0}^{\mathcal{M}}$ in this case.
Next, assume $\lambda>0$. Let us define
$\mathbf{q}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\frac{\lambda b}{u(\mathbf{x})}\cdot\mathbf{p}\,.$
Note that $\mathbf{q}\leq\mathbf{p}$. This follows since $\lambda b=\lambda\left\langle\mathbf{p},\mathbf{x}\right\rangle=\left\langle\mathbf{g},\mathbf{x}\right\rangle\leq u(\mathbf{x})$, where the final inequality is by concavity and $u(\mathbf{0})=0$. Thus,
$g_{j}\leq\frac{u(\mathbf{x})}{b}q_{j}\,,\quad\forall j\in{\mathcal{M}}\,,\,\mbox{with equality if }x_{j}>0\,.$
According to Lemma 4.2, $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
Note that $\mathcal{GD}^{u}(\mathbf{q},b)=\mathcal{GD}^{u}(\mathbf{p},u(\mathbf{x})/\lambda)$. Utilizing Definition 2.14(ii) for $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{p},u(\mathbf{x})/\lambda)$ and $b\leq u(\mathbf{x})/\lambda$, there exists $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $\mathbf{y}\leq\mathbf{x}$ as required. ∎
We next show that for any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$, and any prices $\mathbf{q}\leq\mathbf{p}$, there exist an aggregate Gale demand such that $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ whenever $q_{j}=p_{j}$.
Lemma B.1.
Consider a Fisher market instance such that $u_{i}(\cdot)$ is $\Sigma$-Gale-substitute utilities for all agents. For any competitive equilibrium $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$, for any $\mathbf{q}\in\cap_{i\in{\mathcal{A}}}\operatorname{pdom}(u_{i})$ such that $\mathbf{q}\leq\mathbf{p}$, there exists $\mathbf{y}_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$ and $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ such that $q_{j}=p_{j}$.
@@
By Lemma 5.2, there exists $\bar{\mathbf{y}}_{i}\in\mathcal{GD}_{i}(\mathbf{p},b_{i})$ for every $i\in{\mathcal{A}}$ such that $\bar{\mathbf{y}}_{i}\leq\mathbf{x}_{i}$. Using Definition 2.14(i), there exists $\mathbf{y}_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ and $y_{ij}\leq\bar{y}_{ij}$ if $q_{j}=p_{j}$. Thus, if $q_{j}=p_{j}$ then $\sum_{i\in{\mathcal{A}}}y_{ij}\leq\sum_{i\in{\mathcal{A}}}\bar{y}_{ij}\leq\sum_{i\in{\mathcal{A}}}x_{ij}\leq 1$. ∎
@@
$\varphi(\mathbf{q})\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\sum_{i\in{\mathcal{A}}}\max_{\mathbf{y}_{i}\in\mathbb{R}_{\geq 0}^{{\mathcal{M}}}}\left(b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q},\mathbf{y}_{i}\right\rangle\right)+\sum_{j\in{\mathcal{M}}}q_{j}.$
The properties stated in the next lemma are easy to verify.
Lemma B.2.
@@
    Given $\mathbf{q}\in\operatorname{dom}(\varphi)$, let $\mathbf{y}_{i}\in\mathcal{GD}_{i}(\mathbf{q},b_{i})$ for all $i\in{\mathcal{A}}$. Then, $\mathbf{1}-\sum_{i\in{\mathcal{A}}}\mathbf{y}_{i}\in\partial^{\vee}\varphi(\mathbf{q})$.
@@
Let $((\mathbf{x}_{i})_{i\in{\mathcal{A}}},\mathbf{p})$ be a competitive equilibrium solution. Consider the dual (5) of the Eisenberg-Gale program (EG), i.e.,
@@
We show that this has an optimal solution with $\mathbf{q}\leq\mathbf{p}$. The theorem then follows by Lemma 5.1. Let us take any optimal $\mathbf{q}^{\star}\in\operatorname{dom}(\varphi)$; we are done if $\mathbf{q}^{\star}\leq\mathbf{p}$. Otherwise, let $J\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\star}_{j}>p_{j}\}$, and let us define $\bar{\mathbf{q}}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
[…]
@@
Let us pick an allocation $(\mathbf{y}_{i})_{i\in{\mathcal{A}}}$ for the prices $\bar{\mathbf{q}}$ as guaranteed by Lemma B.1: $\sum_{i\in{\mathcal{A}}}y_{ij}\leq 1$ for all $j\in{\mathcal{M}}$ with $\bar{q}_{j}=p_{j}$; in particular, for all $j\in J$. According to Lemma B.2(ii), we get a subgradient $\bar{\mathbf{g}}\in\partial^{\vee}\varphi(\bar{\mathbf{q}})$ defined as $\bar{\mathbf{g}}\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\mathbf{1}-\sum_{i\in{\mathcal{A}}}\mathbf{y}_{i}$; this has the property $\bar{g}_{j}\geq 0$ for all $j\in J$. We then have
@@
It is left to show that $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)$. Let $K\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}\{j\in{\mathcal{M}}\,:\,q^{\star}_{j}=0<p_{j}\}$. Clearly, $K\cap J=\emptyset$. If $K=\emptyset$, then Lemma A.1(iii) already implies $\bar{\mathbf{q}}\in\operatorname{pdom}(u)$, since $\mathbf{p}\geq\bar{\mathbf{q}}$ and $\operatorname{supp}(\bar{\mathbf{q}})=\operatorname{supp}(\mathbf{p})$. Otherwise, for $k=1,2,\ldots$, let us define $\mathbf{q}^{(k)}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ as
[…]
@@
We have $\mathbf{q}^{(k)}\in\operatorname{pdom}(u)$ again by Lemma A.1(iii). As in the first part, we take a subgradient $\mathbf{g}^{(k)}$ as guaranteed by Lemma B.2(ii), with the properties that $g^{(k)}_{j}\geq 0$ for all $j\in J$ and also $\mathbf{g}^{(k)}\leq\mathbf{1}$. We can write
@@
$\varphi(\mathbf{q}^{\star})\geq\varphi(\mathbf{q}^{(k)})+\left\langle\mathbf{g}^{(k)},\mathbf{q}^{\star}-\mathbf{q}^{(k)}\right\rangle=\varphi(\mathbf{q}^{(k)})+\sum_{j\in J}g^{(k)}_{j}(q^{\star}_{j}-q^{(k)}_{j})-\sum_{j\in K}g^{(k)}_{j}q^{(k)}_{j}\geq\varphi(\mathbf{q}^{(k)})-\frac{\|\mathbf{p}\|_{\infty}}{k}|K|\,.$
@@
Here, we use in the first equality that $q^{\star}_{j}=q^{(k)}_{j}$ for $j\notin J\cup K$, and $q^{\star}_{j}=0$ for $j\in K$. The last inequality uses $g^{(k)}_{j}\leq 1$ and the definition $q^{(k)}_{j}=p_{j}/k$. Thus, Lemma A.1(v) is applicable with $M=\varphi(\mathbf{q}^{\star})+|K|\cdot\|\mathbf{p}\|_{\infty}$, using that for each $j\in{\mathcal{M}}$, $\max_{\mathbf{y}_{i}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}b_{i}\log u_{i}(\mathbf{y}_{i})-\left\langle\mathbf{q}^{(k)},\mathbf{y}_{i}\right\rangle\leq\varphi(\mathbf{q}^{(k)})$. This shows that $\bar{\mathbf{q}}\in\operatorname{pdom}(u_{i})$ for all $i\in{\mathcal{M}}$, and therefore $\bar{\mathbf{q}}\in\operatorname{dom}(\varphi)=\cap_{i\in{\mathcal{M}}}\operatorname{pdom}(u_{i})$. ∎
Appendix C Separable utilities are $\Sigma$-Gale-substitutes